Solve each of the systems by addition. If a unique solution does not exist, state whether the system is inconsistent or dependent.
x + 5y = 10
-2x - 10y = -20
Thanks for any help you can provide.

Answer by stanbon(75887)

1st: x + 5y = 10
2nd: -2x - 10y = -20
---------------
Multiply 1st be 2 to get:
3rd: 2x+10y=20
Add 2nd and 3rd to get:
0=0
----------
Conclusion: the system is dependent.
Cheers,
Stan H.